Journal of Lie Theory Volume 6 (1996) 179–190 ◯^{C} 1996 Heldermann Verlag

Torus actions on compact quotients

Anton Deitmar

Communicated by K.-H. Neeb

Abstract. A Lefschetz formula for actions of noncompact tori on compact quotients of Lie groups is given.

Math. Inst. INF 288 69126 Heidelberg Germany

Received March 21, 1996 and in final form June 13, 1996

ISSN 0949–5932 / $2.50 ◯^{C} Heldermann Verlag